### v2

# A shared compilation stack for distributed-memory parallelism in stencil DSLs

Abstract.
Domain Specific Languages (DSLs) increase programmer productivity and provide high performance. Their targeted abstractions allow scientists to express problems at a high level, providing rich details that optimizing compilers can exploit to target current- and next-generation supercomputers. The convenience and performance of DSLs come with significant development and maintenance costs. The siloed design of DSL compilers and the resulting inability to benefit from shared infrastructure cause uncertainties around longevity and the adoption of DSLs at scale. By tailoring the broadly-adopted MLIR compiler framework to HPC, we bring the same synergies that the machine learning community already exploits across their DSLs (e.g. Tensorflow, PyTorch) to the finite-difference stencil HPC community. We introduce new HPC-specific abstractions for message passing targeting distributed stencil computations. We demonstrate the sharing of common components across three distinct HPC stencil-DSL compilers: Devito, PSyclone, and the Open Earth Compiler, showing that our framework generates high-performance executables based upon a shared compiler ecosystem.
1. Introduction
Most DSLs are purposefully designed for specific problem domains, providing focused and specialized language constructs tailored to those domains. They capture concepts for a specific domain, yet a large portion of their code base is dedicated to reasoning about generic concepts in HPC. These include the generation of directives for shared-memory parallelism, message-passing communications for distributed-memory parallelism, vectorization, arithmetic (factorization, sub-expression elimination), and loop optimizations (blocking, fusion, fission). These general-purpose optimizations are often combined with domain-specific ones. After lowering away domain-specific knowledge, compiling an intermediate representation (IR) to executable code often requires the DSL library developers to implement standard compiler passes. Sharing the developed technology between distinct projects is challenging due to integration issues with different programming languages, lack of standardization, and maintenance support.
MLIR (Lattner et al., 2021) has partially solved this interoperability problem, especially for deep-learning libraries, by providing an extensible and community-driven unified IR. However, efforts are still needed to bridge the gap between HPC concepts and MLIR. This work aims to address this gap by introducing key HPC abstractions designed for interoperability with existing MLIR dialects. We utilize xDSL (Fehr et al., 2023), a Python-native clone of the MLIR toolchain to achieve this goal. In this work we focus on explicit finite-difference (FD) stencil computations on structured grids. While FD stencils serve as a meaningful and illustrative example to showcase the applicability of the proposed techniques, the intention is to contribute to the advancement of HPC practices more broadly, with the FD stencil computation acting as a representative case study within a larger ecosystem of potential applications. This paper makes the following contributions:
A set of HPC-specific compilation components comprising SSA dialects and lowerings between them:
An SSA dialect to facilitate automated domain decomposition for distributed-memory execution of stencil kernels via message-passing (section 4.2).
An SSA dialect for message passing as a set of modular operations in a standardized SSA-based IR (section 4.3);
A prototype implementation of abstraction-sharing compilation stack for two HPC stencil-DSL compilers, PSyclone and Devito, based on the concepts of SSA and Region, utilizing the MLIR and xDSL compiler frameworks and extending concepts from the Open Earth Compiler (section 5).
A performance evaluation demonstrating that our approach offers competitive performance for a range of FD stencil computations, compared to the existing domain-specific compiler stacks, for CPU shared- and distributed-memory parallelism, GPUs and FPGAs running at scale on the ARCHER2 and Cirrus supercomputers (section 6).
The paper is organized as follows: section 2 discusses our compiler-centric approach, section 3 introduces MLIR for abstraction sharing, section 4 presents our SSA dialect abstractions that are the main contributions of this work, section 5 presents a prototype implementation for the stencil DSLs used in this work to validate our contributions, and section 6 evaluates the performance of our final software stack. Sections 7 and 8 refer to related work and conclude with future directions.
2. Stencil DSLs with Shared Abstractions
We propose a compiler-centric approach where stencil DSLs maximize the sharing of their implementation (e.g. common abstractions from their application domains or commonly used HPC abstractions) while maintaining the community-tailored interfaces and workflows that have proven critical to communicate effectively with their domain experts. By compiler-centric we mean that we build and contribute compiler infrastructure and leverage capabilities and optimizations provided by compilation frameworks, such as transformation passes, abstractions, performance optimizations, portability, and productivity.
We focus on three DSLs targetting explicit FD stencils on structured meshes, Devito (Luporini et al., 2020), the Open Earth Compiler (Gysi et al., 2021), and PSyclone (Adams et al., 2019).
FD-stencils computation pattern is an iterative computation update of an element in an -dimensional spatial grid at time as a function of neighboring grid elements (space dependencies) at previous timesteps (time dependencies). Figure 2 illustrates a 1D-3pt Jaocobi stencil and its flow dependences. Each point is updated using values from the previous timestep and the right and left neighbors. Arrows illustrate the flow dependencies. Halo points (grey) are read-only and extend the computational domain by the size of the stencil radius.
All the prementioned DSLs and their compiler implementations are developed as siloed, standalone frameworks with zero code sharing (fig. 1(a)). However, despite this lack of code sharing, all three stencil compilers use similar concepts and transformations for domain-specific and general-purpose HPC optimizations. To address this lack of code sharing, we propose a novel shared compilation stack for these three frameworks (fig. 1(b)), where domain-specific optimizations and HPC concepts are shared. User interfaces differ, but the underlying computational abstractions reason about common concepts.
Devito offers a DSL and compiler framework to automate the FD method and is primarily used in seismic inversion and medical imaging. The Open Earth Compiler was built for weather and climate models using stencils. Finally, PSyclone is designed for weather, climate, and ocean models. Devito is an embedded DSL and compiler framework written in Python. The Open Earth compiler is a domain-specific compiler offering a frontend at the stencil specification level, and PSyclone uses Fortran, augmented with specific coding conventions to connect with the HPC community.
Nevertheless, all three frameworks eventually arrive at the domain-specific concept of stencil computations. Stencil kernels (and sequences thereof) are optimized and translated into imperative code that runs on HPC machines. Hence, both the domain-specific optimization at the stencil level and the code generation for an HPC machine are conceptually similar.
We show that these similarities can be exploited to reuse and share the lower-level compilation flow across the three frameworks. The key to enabling such reuse is to share abstractions across the stencil DSL compilers. The compiler community has embraced the concepts of static single assignment (SSA) and nested control flow regions as recommended program expression and transformation approaches. The MLIR project (Lattner et al., 2021) has recently demonstrated that these concepts can be instantiated for different domains (see section 3) and used to provide a principled way to exchange information through different stages of the compilation pipeline. We embrace these concepts to design a shared compilation flow (fig. 1(b)) that connects these three stencil DSL compilers. Specifically, we demonstrate that all three compilers can use a common stencil abstraction to optimize and generate code for their stencil kernels. Starting from this stencil-level abstraction (which is generated from disparate, domain-specific inputs), the shared infrastructure translates the common stencil abstractions to stencil kernels and leverages established SSA-based compiler IRs for loops, arithmetic, and memory operations. For building a stencil-level abstraction we adopt the stencil and GPU IRs in the form of MLIR dialects111https://mlir.llvm.org/docs/Dialects/. These two dialects, introduced in Gysi et al. (2021), are discussed in detail in section 4 where we expand their features to better support the stencil DSLs studied here. This was selected as a least-friction way to lower stencil-related concepts to MLIR and LLVM. The Open Earth Compiler (Gysi et al., 2021) has already successfully targeted the lowering of stencil concepts to the MLIR framework. We reuse and extend its infrastructure, and more specifically, its Stencil IR (fig. 1(a)), used as the basis for the stencil dialect in this work (refer to section 4.1 for more details).
This lowering process involves a series of passes across SSA-based compiler IRs. Additionally, we introduce new IRs based on the SSA+Regions abstraction (section 3) for lowering the stencil dialect IR to distributed memory parallelism via the Message Passing Interface (MPI) standard. The result is a comprehensive lowering stack that automatically generates shared- and distributed-memory and GPU-accelerated computations for Devito, the Open Earth Compiler, and PSyclone.
3. Sharing Abstractions through IRs
The abstractions transformed by compilers must concretely be realized in code as some IR. IRs that maintain SSA form (Cytron et al., 1991) have proven particularly useful as their direct encoding of a program’s data flow allows transformations without requiring complex analyses. Recently, the MLIR project (Lattner et al., 2021) demonstrated that adding hierarchical structure to SSA-based IRs via nested regions facilitates the modeling of higher-level domain-specific constructs, enabling development of compilers in deep learning (Paszke et al., 2019; Abadi et al., 2016, 2015), fully-homomorphic encryption (Govindarajan and Moses, 2020), and hardware design (Eldridge et al., 2021; Majumder and Bondhugula, 2024). This SSA+Regions concept provides a principled way to interface distinct domain-specific IRs, which we adopt as a foundational principle in our compiler stack.
[…]
While various compiler frameworks use SSA-based IRs, the introduction and use of regions is more recent and central in MLIR and xDSL. MLIR is an LLVM subproject that supports SSA and regions in a performance-focused C++ compiler framework. It comes with pre-defined abstractions for arithmetic, structured control flow, memory references, GPUs, and many more. The xDSL project is a Python-native clone of MLIR that provides an embedding of SSA and regions into Python facilitating interoperability with Python-based DSLs. Similar to MLIR, xDSL is open-source and publicly available on GitHub222https://github.com/xdslproject/xdsl. Both frameworks are built on the previously introduced concepts of SSA and Region and share the same textual representation to share infrastructure without tight coupling of code. Users may introduce their own abstractions by defining new dialects specific to their application domain. Since most DSLs we are targeting are written in Python, we start our compilation pipeline from xDSL and define our custom abstractions there. We develop dialect-specific lowerings from xDSL to existing MLIR dialects and then hand off to the MLIR toolchain where we apply further lowerings to target particular hardware.
4. SSA Compiler Abstractions for HPC
The memory requirements of computational science problems in HPC often exceed the capacity of DRAM on a single platform. Fortunately, many HPC problems lend themselves to spatial domain decomposition, allowing for the distribution of computational tasks across multiple compute nodes within a cluster. Thus, domain decomposition and message-passing abstractions are considered integral parts of HPC.
While MLIR provides several dialects that also benefit HPC users, such as omp for OpenMP shared-memory parallelism (SMP), gpu for generic GPU programming across NVIDIA and AMD GPUs, and vector for vectorizing computations, it does not provide a means to control distributed memory parallelism (DMP).
We extend the capabilities of the stencil dialect introduced in Gysi et al. (2021) to target multi-node HPC clusters through a novel, extended version of this dialect (section 4.1). This dialect captures the geometric decomposition pattern of where the domain is decomposed into rank-local domains. The rank-local domains communicate at predefined points in the computation. One of the uses of the dmp dialect is in the developed lowerings from the existing stencil dialect. While the stencil dialect contains the mathematical abstract representation of the problem, the dmp carries a more concrete realization. Although still at a fairly high level from a parallelism perspective, we consequently lower to the mpi dialect, an even lower-level realization of the parallelism and decomposition.
Frameworks can represent their stencil computations using the stencil dialect, benefiting from common IR and semantics. The dmp dialect is an abstraction for distributed-memory parallel (DMP) operations, such as data exchanges through message passing (section 4.2). Through the dmp dialect, frameworks that have expressed their IR in stencil benefit from automatically decomposing and distributing the computational domain to the MPI ranks that are used for the computation. We refer to the data owned by an MPI rank as “rank-local” data. Finally, we introduce an MPI-based dialect (section 4.3) that mirrors MPI’s point-to-point and collective communications and serves as a lowering target for the dmp dialect. Communications are expressed using this novel dialect, facilitating the development of codes leveraging message passing. The implementation of stencil, dmp, and mpi dialects is open-source and available online.
4.1. A Stencil Dialect
The initial implementation of the stencil dialect was introduced as part of the Open Earth compiler (Gysi et al., 2021), to represent stencil computations and their sequences by capturing essential information, including iteration space bounds and stencil update patterns. This dialect acts as an intermediate step, helping to close the semantic gap between the scientist’s problem description and the concrete implementation on CPUs and GPUs.
A major benefit of the stencil dialect is that it is problem, domain, and hardware independent, in contrast, for instance, to other approaches such as Essadki et al. (2023) and Li et al. (2024). For example, Essadki et al. (2023) introduced a cfd MLIR dialect for stencils associated with Computational Fluid Dynamics (CFD) problems, with operations and transformations focused on this specific problem. Consequently, their dialect is tied to one class of DSL rather than providing the ability to generalize across domains. By contrast, the work described in Li et al. (2024) targets the porting of stencil applications to the Sunway SW26010 heterogeneous many-core processor, with the dialect containing operations influenced by the target machine architecture. Instead, when using the stencil dialect from Gysi et al. (2021), a programmer’s code is lowered to a general-purpose description of a stencil computation independent of the hardware architectures. A major contribution of Gysi et al. (2021) was to demonstrate that this general description of a stencil computation contains enough information for subsequent transformations to effectively operate upon when generating optimized HPC target executables.
In this work, we integrated the stencil dialect into the xDSL framework and extended the dialect to enable integration with our DMP and MPI abstractions, facilitating support for various lowering targets, including CPUs, GPUs, and FPGAs. Due to its general purpose nature, the stencil dialect contains 11 operations in total, compared, for instance, to Essadki et al. (2023), which only contains two. The following list briefly highlights the most important operations and types in the stencil dialect.
[…]
stencil.field is a type that represents the memory buffer from which stencil input values will be loaded, or to which stencil output values will be stored
[…]
stencil.store is an operation that writes values to a field on a user-defined range.
[…]
stencil.load is an operation that takes a field and returns its values.
[…]
Section 4.1 provides an example of the stencil dialect to compute a 1-dimensional 3-point Jacobi calculation (see fig. 2). {listing}[bp] {mlir} -> !temp<?xf64> -> !temp<?xf64> // stencil.return
[…]
Working with authors of the stencil dialect, we have made various enhancements. For instance, in the original dialect, bounds information was encoded as attributes of MLIR operations. However, this required additional analysis and transformations in lowering from the logical coordinates of the stencil representation to zero-based memory operation indexing. To avoid these complexities, we modified the dialect to associate domain bounds information with the values themselves through their types. This allows any operation using stencil-related types to access this information directly through their operands. It not only simplifies the existing stencil transformations but enables independent lowering to the DMP abstraction, decoupling this from stencil-specific operations that define logical coordinates.
Previously, the stencil dialect was only capable of representing 3D stencils and we have also enhanced the dialect so that it can be used with stencils of any number of dimensions. Not only does this enable us to run simpler, 1D and 2D problems, but furthermore provides flexibility for more complicated future domains. The original dialect was specifically tailored to target GPUs and-so we have enhanced the stencil transformations by providing an additional lowering pipeline which is better suited for shared memory parallelism by leveraging loop tiling to improve data locality.
[…]
Limitations
The main functional limitation of the stencil dialect is the requirement for compile-time known bounds. Whilst this requires recompilation of code when the problem sizes changes, the performance benefits delivered by having this information statically available during compilation are significant. For example, known bounds enable constant-folding of most of the memory access address computations and thus reduce register pressure which is critical to increasing parallelism on GPUs. Due to the nature of DSL compilers, propagating bounds from the call site to the stencil expression is usually not difficult in practice, while the existence of compile-time known bounds greatly aids in implementing more complex compiler analysis and optimization passes.
Another limitation is the absence of a high-level representation for modeling boundary conditions, and this feature is scheduled for future work. The current version of the dialect is flexible enough to manually encode boundary conditions as conditionals over the memory accesses. However, expressing these at a higher level could lead to more opportunities for optimizing code generation and enhancing the supported use cases for the users.
4.2. dmp dialect: An IR for Domain Decomposition
This section introduces a novel dialect for distributed memory parallelism (DMP), the dmp dialect. dmp is used to express parallel communication patterns as modular building blocks that lower computational dialects, such as stencil to an IR for DMP. dmp offers a mechanism for describing the exchange of rectangular subsections of data among nodes. Additionally, we introduce a pass to convert a stencil kernel into an IR for distributed computations. This pass allows for optimizing intra- and inter-node performance gains using the dmp dialect for targeting large-scale runs.
[…]
The design of the dmp dialect is directly influenced by design choices made in the stencil dialect. The compile-time known bounds of the stencil expression greatly aided in reducing the complexity of the dmp operations and transformations, demonstrating that the simplicity of the stencil dialect directly benefits the available compiler optimizations.
Limitations
The dmp dialect is a prototype implementation of domain decompositions for stencil computations. It is designed to operate on contiguous, hyper-rectangular subsections of the domain and, therefore, does not support operations on sparse or non-rectangular-shaped domains. Extending the dialect to cover more data layouts is projected for future work.
Another limitation arises due to the fact that one exchange operation corresponds to one halo exchange, making the implementation of optimizations such as communication-computation overlap and exchanging multiple halos simultaneously more complicated than it needs to be. Finding a suitable extension to the dmp dialect for more application domains and more efficient rewrites is also projected as an interesting direction for future work.
4.3. mpi dialect: An IR for Message Passing
Our final contribution to the MLIR ecosystem is the MPI dialect, designed to mirror MPI’s point-to-point and collective communications. The mpi dialect has already been upstreamed to the MLIR compiler toolchain333https://mlir.llvm.org/docs/Dialects/MPI/. This dialect serves as a target for our dmp dialect and is lowered to MPI library function calls.
Our work here comprises the dialect itself, containing the abstract operations and types. The operations correspond to the MPI calls, while the types represent MPI types such as request handles, communicators, and data types. Additionally, we have introduced operations to reduce the friction between the MPI and the MLIR ecosystems. These operations enable the expression of common MPI concepts outside of the library calls, such as request object lists and interactions with memory references. Because MLIR and the LLVM backends lack an intrinsic notion of MPI, we have also developed a lowering that will convert the MPI dialect into the underlying function calls and associated memory allocations using the low-level llvm and func MLIR dialects. The contributed mpi dialect currently supports the following subset of the MPI 1.0 standard operations:
Blocking and Non-blocking Point-to-Point communications: MPI_Send, MPI_Recv, MPI_Isend, MPI_Irecv
Nonblocking request operations: MPI_Test, MPI_Wait and MPI_Waitall
Blocking reductions: MPI_Reduce and MPI_Allreduce
Broadcast and Gather collective: MPI_Bcast, MPI_Gather
Process Management: MPI_Init, MPI_Finalize, MPI_Comm_rank and MPI_Comm_size
While this subset of MPI operations provides the necessary building blocks for the current needs of our MPI applications, the design allows for extensions. Additionally, we provide operations that simplify targeting mpi from higher levels of abstraction. Our compiler transformations and lowerings leverage these operations to implement DMP.
One of the lowerings we provide transforms dmp to mpi operations. Note that there is nothing MPI-specific in the dmp dialect, allowing targeting other communication libraries if desired. Lowering to mpi involves several steps, including allocating temporary buffers, building the MPI exchange mapping, packing/unpacking data to/from buffers, and issuing non-blocking send/receive calls. Figure 4 shows the resulting transformation, with some details omitted for readability, such as setting skipped request objects to the null request and temporary send/receive buffer allocations and de-allocations. Since MPI communication often happens inside loops, any loop invariant calls are hoisted as part of this transformation.
As LLVM has no concept of MPI, we lower these operations to regular function calls using the func dialect. Motivated by the distribution on our target platform (ARCHER2), we support the mpich implementation of the MPI standard. To make use of MPI, it is usually required to include the implementations C header file, a notion not supported by MLIR. Instead, we extract magic values from our library’s header file and substitute them for e.g. datatype constants during the lowering process. This makes our provided MPI lowering specific to the mpich library, but this strategy can be adapted to other MPI libraries like OpenMPI. Futhermore, there is an ongoing effort to provide one standard MPI ABI, greatly simplifying this problem (Hammond et al., 2023).
[…]
5. Lowering stencil DSLs to MLIR
This section briefly introduces the frameworks used and then details the methodology to leverage the synergies exploited from the contributed shared infrastructure. Figure 1(b) schematically presented the shared infrastructure contributed by this work. Devito and PSyclone lower their IRs to the stencil, dmp and mpi dialects, benefiting from the advantages presented in section 4. Additionally, they leverage “out-of-the-box” lowerings to other dialects optimized for HPC, such as omp, gpu designed for parallel execution on CPUs and GPUs.
5.1. Devito
Devito (Luporini et al., 2020; Louboutin et al., 2019) is an open-source Python DSL and compiler framework widely used in academia and industry, aiming to ease the development of HPC finite-difference PDE solvers. Devito’s symbolic DSL is based on Sympy (Meurer et al., 2017), facilitating expressing and solving PDE simulations. Devito’s compiler framework automates the conversion into optimized FD kernels for efficient execution on CPUs and GPUs. An example of the Devito DSL for modeling a 2D heat diffusion problem is shown in section 5.1. {listing}[htbp] {python} # Model the problem and automatically generate code grid = Grid(shape=(126,)) u = TimeFunction(name=’u’, grid=grid, space_order=2) eqn = Eq(u.dt, 0.5 * u.laplace) op = Operator(Eq(u.forward, solve(eqn, u.forward))) # JIT-compile the code and run op(t=timesteps) The high-level symbolic code to model 1-dimensional heat diffusion in Devito DSL. Users can only focus on modelling, all machinery for optimized HPC code generation is abstracted away.
Devito’s separation of concerns is benefiting interdisciplinary research and development by reducing time-consuming and error-prone practices widely encountered in stencil codes. Devito uses a high-level symbolic mathematic-textbook-like abstraction as input that is very close to what mathematicians, physicists and geoscientists are familiar with. It assists interdisciplinary scientists with the rapid development of simulation codes for wave propagators or other PDE-dominated problems, as it facilitates the development of mathematically sophisticated PDE simulations, the implementation of stencil kernel adjoints, boundary conditions, sparse non-aligned-to-the-structured-grid operations (Bisbas et al., 2021) and others. All the compiler optimizations are abstracted away from the DSL user: arithmetic and loop performance optimizations (e.g., common sub-expression elimination, loop tiling, factorization), and backend-specific optimizations, such as SIMD vectorization, SMP and DMP, and GPUs. Devito’s compiler automates optimization passes, allowing users to apply further tuning heuristically. On completion of the compiler passes, high-performing C code for efficient PDE solvers is automatically generated.
The shared Devito-PSyclone approach to lowering stencil IRs to xDSL is shown in fig. 6. The upper part of the fig. 6 illustrates an overview of lowering Devito’s stencil representation to xDSL and, subsequently, MLIR and LLVM. Initially, Devito’s symbolic input undergoes lowering to a representation, which encapsulates the FD-stencil representation regarding iteration spaces, memory accesses, mathematical operations, and many more.
The primary Devito object that contains the necessary semantics in Devito is a group of expressions. Subsequently, these are parsed and lowered to stencil expressions (fig. 5). Building upon the example presented in section 5.1, fig. 5 shows some of the IR that Devito is using to represent mathematical equations. We parse info on read and write accesses from Devito’s IR, and use this information to construct expressions using the stencil dialect, which will then be further lowered as described in fig. 4.
While a number of dialects are used when lowering Devito to the xDSL representation (refer to fig. 6), some of the main corresponding matches involved while visiting Devito structures are: the translation of mathematical expressions using operations from the func and arith, translation of Devito’s memory accesses to memref and stencil dialect, translation of the problem dimensions to structured control flow using the scf dialect and targetting specific backends using the mpi and gpu dialects. The integration with Devito happens at the highest possible level of Devito’s compiler IRs to evaluate the optimization capabilities MLIR offers compared to the main Devito optimization pipeline. This approach helps to compare pipelines that do not share any optimizations and start from the same level of abstraction.
The contributed stencil dialect facilitates the lowering from Devito to build the FD structured grid, the necessary ghost-cell read-only region, the stencil coefficients, and memory accesses. Subsequently, we add the temporal and spatial loops, including time-buffering. We benefit from applying transformation and optimization passes from the shared infrastructure available from the MLIR community444https://mlir.llvm.org/docs/Passes/ such as cse, loop-invariant-code-motion, fold-memref-alias-ops and convert-scf-to-openmp. Finally, HPC code leveraging SMP, DMP for CPUs, and GPUS is generated for several targets and then compiled and executed. We evaluate these variants against standalone Devito in section 6.
5.2. PSyclone
PSyclone (Adams et al., 2019) is a DSL for Fortran codes, enabling scientists to write their kernels in Fortran, a language that they are already familiar with, and then leverage a translation layer that abstracts the mechanics of the computation and parallelism. PSyclone is popular with weather and climate, for instance, the Met Office’s next-generation weather model, LFric (Melvin et al., 2017), is using this technology and they are also involved in several other activities around PSyclone such as the NEMO ocean model port (Porter et al., 2018). The major design feature of PSyclone is to enable a separation of concerns, where scientists develop the algorithm and kernel layers, both representing the mathematics of their problem and directing the sequence of maths operations that should be run, with the tool then generating the PSy layer. This PSy layer connects the algorithm and kernel and layers together, for instance, applying distributed memory parallelism. Furthermore, the PSyclone compiler will analyze the scientist’s provided code and undertake optimizations and transformations, such as applying OpenMP or OpenACC at the loop level for threaded parallelism and GPU acceleration respectively.
However, the compilation stack of PSyclone is siloed with a bespoke translation layer that has been written in Python and, after parsing the Fortran source code, builds an internal intermediate representation based in a Directed Acyclical Graph (DAG) form. This IR is then operated upon by transformations bespoke to the PSyclone compiler before the IR is transformed into either Fortran or C. Domain-specific APIs enhance the core functionality of PSyclone, where an API contains a specialized set of transformations that target individual codes or scientific areas. For the results discussed in this work we leverage the NEMO API which has been developed to enable PSyclone to target the NEMO ocean model (Porter et al., 2018).
5.2.1. Lowering PSyclone to xDSL/MLIR/LLVM
Figure 6 illustrates the shared Devito-PSyclone approach for lowering abstractions to xDSL, then building upon the shared transformations, existing MLIR dialects and new HPC dialects presented in this paper. Consequently, the PSyclone DSL is now a thin abstraction layer a-top a common shared ecosystem which is also used by Devito as described in section 5.1. We still use the lexing and parsing of PSyclone to build the PSyclone IR. This is then passed directly to our PSyclone xDSL backend to generate our own PSy IR, an xDSL-based IR that closely resembles PSyclone’s own IR. Whilst this is still heavily PSyclone dependent, as it is in DAG form, there is a rich structure to the representation that appropriate transformations can exploit and a one-to-one mapping between PSyclone’s existing IR and our xDSL PSy-IR.
An example of such a transformation that can be applied at this stage by the PSyclone xDSL backend is the identification of stencils from Fortran loops. Any identified stencils are represented in the PSy-IR dialect which is then lowered to SSA form. Once in SSA form the flow is within the common xDSL ecosystem and there are two main dialects that PSy-IR is lowered to. The stencil dialect, with the flow exactly as described for Devito in section 5.1; and the FIR dialect, used by Flang to represent Fortran. Lowering the non-stencil aspects into this standard Fortran dialect not only means that we preserve the semantics of the surrounding code but also the escape hatch of PSyclone which is the ability to handle all Fortran that is not transformed by PSyclone itself.
Whilst a mix of dialects is natural when working with MLIR, this raises a challenge with the LLVM tooling because the FIR dialect is separated from other dialects. Flang, which defines the FIR dialect, is only capable of working with that dialect and a very small subset of standard dialects such as arith and func. Likewise, mlir-opt, which is used to operate on the standard dialects is unaware of FIR. These different standard dialects are crucial; for instance, the gpu dialect is required for launching on the GPU and the memref dialect for handling memory within stencils, but Flang is not able to process this.
As shown in fig. 6, the FIR and stencil aspects are kept separate, with Flang used to build the LLVM-IR from the FIR dialect and mlir-opt the lowered dialects from transformations on the stencil dialect which are wrapped in functions for each stencil. Ultimately, this generated LLVM-IR is passed to the corresponding LLVM backend for the target architecture, and compiled into object files that are linked together to form the target executable where FIR can then call into the stencil dialect components by issuing the appropriate function calls.
6. Evaluation
This section presents the experimental evaluation of this paper’s contributions versus the standalone Devito and PSyclone compiler stacks. The systems used for the experiments were: (a) ARCHER2 HPE Cray EX Supercomputer nodes comprising a dual AMD EPYC 7742 64-core 2.25GHz processor with 128 cores. Each node has 8 NUMA regions and 16 cores per NUMA region, 32kB of L1, 512kB of L2 cache per core, 16MB L3 cache for every 4 cores and supports AVX2 vectorization. It has an HPE Slingshot interconnect with 200 Gb/s bandwidth, dragonfly topology, and nodes are organized into groups of 128. CCE’s Cray Clang 11.0.4 was used. (b) Cirrus GPU compute nodes consisting of four NVIDIA Tesla V100-SXM2-16GB (Volta) GPU accelerators. On Cirrus, we used version Cray Clang 11.4.100, nvc 22.11 from NVIDIA HPC SDK 22.11, and CUDA 11.6.
We use modified versions of xDSL, PSyclone, and Devito. Experiments were executed with the Devito code available here [xdslproject/devito] and xDSL code available here [xdslproject/xdsl]. For the performance evaluation of the stencil kernels we use GPts/s (a.k.a GCells/s) for throughput.
6.1. Devito
We use two benchmarks from CFD and seismic imaging: (i) heat diffusion, a Jacobi-like stencil, and (ii) the isotropic acoustic wave equation. We benchmark both problems in 2D and 3D for varying space discretization orders (SDO) of 2, 4, and 8, resulting in 5-, 9-, 13-pt for 2D and 7-, 13-, 19pt stencils for 3D. The isotropic acoustic wave equation uses a 2nd order accurate approximation in time, resulting in more points being read at the time dimension. On ARCHER2, the problem size is for the 2D and for the 3D case. On Cirrus, the problem size is for the 2D and for the 3D case. The simulation time in timesteps is for the 2D and for the 3D case. For the strong scaling experiments, we focus on the 3D problems with a bigger working set. For all benchmarks, we use a 32-bit single-precision floating point numbers.
Figure 7 evaluates xDSL-Devito versus the native Devito. This experiment uses 8 MPI ranks and 16 OpenMP threads per rank on an AMD EPYC 7742 node of ARCHER2, to better suit the NUMA architecture of the underlying architecture. The Devito benchmark was optimized using the full suite of flop reduction, DMP, SMP, and SIMD vectorization optimizations (Luporini et al., 2020; Bisbas et al., 2023). Devito’s solvers are highly efficient, as established on roofline plots in related work (Luporini et al., 2020; Bisbas et al., 2021), ensuring we compare against a highly competitive baseline. xDSL-Devito’s performance outperforms Devito for low arithmetic intensity (AI) kernels. However, Devito’s flop reduction optimizations pay off for higher AI. One reason for this is the limited vectorization performance of our current lowered LLVM IR, which is essential for stencil kernels (Henretty et al., 2011).
Figures 8(a) and 8(b) show the strong scaling of the heat diffusion and acoustic wave stencil kernel, respectively. We benchmark both kernels for a 3D problem and SDO of 4. We use a whole ARCHER2 group of 128 nodes scaling up to 1024 MPI ranks, totaling 16384 cores. We observe that xDSL-Devito exhibits strong scaling that may not match Devito’s performance but still maintains the scaling trend. This regression is expected, as Devito supports more advanced communication techniques, consisting of 3D diagonal exchanges leading to more robust and efficient scaling. More details on the domain decomposition techniques employed by Devito can be found in Bisbas et al. (2023). Extending computation/communication patterns in the dmp dialect is planned for future work. There is definitely space for more improvements on xDSL’s end, which is discussed in section 8.
Figure 9 presents the GPU evaluation. We run on a V100, using OpenACC GPU offloading for Devito555DevitoPRO’s license would be needed to compare against CUDA, reporting the best performance among using collapse(2), collapse(3) or tile(32,4,8) for the spatial loops of the stencil. xDSL-Devito’s kernels use MLIR lowerings to CUDA (Gysi et al., 2021) and use tiled execution. These lowerings allow us to benefit from an out of the box CUDA backend. Evaluation shows, as expected, that Devito/xDSL CUDA kernels outperform Devito’s OpenACC for the higher-intensity wave kernels while being mostly on par with the rest of them. Profiling using nsys666https://developer.nvidia.com/nsight-systems, we observe that the main reason is superfluous synchronization overhead on each kernel launch, which is amortized in the larger kernels by the kernel runtime itself. The MLIR version used does not yet offer a built-in solution for this.
6.2. PSyclone
We have selected two benchmarks to explore performance for our approach with PSyclone, the first is the Piacsek and Williams advection scheme (Piacsek and Williams, 1970), commonly used by Met Office codes such as the MONC high-resolution atmospheric model (Brown et al., 2015), for calculating the movement of quantities through the atmosphere due to kinetic effects (e.g. wind). The second benchmark is the tracer advection kernel from the NEMO ocean model and part of the PSyclone benchmark suite (Science and Technology Facilities Council (2021), STFC). Both these schemes are extracted from production codes and the benchmarks are different in a couple of crucial ways. Firstly, PW advection contains three separate stencil computations across three fields, whereas tracer advection comprises 24 stencil computations across six fields. Each stencil computation involves many individual calculations, and for tracer advection, there are many more computations than fields due to dependencies between computations generating intermediate results. Furthermore, the tracer advection benchmark contains an outer loop where the entire stencil calculation is executed for 100 iterations.
Figure 10(a) provides a performance comparison of using our approach, xDSL-PSyclone, against PSyclone where the target code is compiled with the Cray compiler, Cray-PSyclone and the GNU compiler GNU-PSyclone. We report throughput across our benchmarks (pw for PW advection and traadv for tracer advection) at different problem sizes. It can be seen that for the PW advection benchmarks the performance delivered by xDSL slightly exceeds that of PSyclone with the Cray compilers, whereas PSyclone with the GNU compiler is performing considerably worse. This demonstrates that the Cray compiler is undertaking numerous HPC optimizations over and above the GNU compiler, and given that xDSL is slightly outperforming the Cray compiler our stack can undertake similar levels of optimizations.
However, in fig. 10(a) it can be seen that the performance is different for tracer advection, where for smaller problem sizes xDSL is considerably slower than PSyclone with both the Cray and GNU compilers, although the performance gap narrows for larger problems. The reason for this is the number of individual stencils, where for the PW advection benchmark the three stencil computations are fused into one single stencil region by xDSL, but with tracer advection there are 18 individual stencil regions due to dependencies. Limitations in the lowering from scf.parallel to the OpenMP dialect result in each stencil region being lowered by MLIR into a separate parallel region. Consequently, the kmp_wait_template function, which spins on a barrier waiting for parallel worker threads to complete, was the most runtime-intensive function when profiling smaller domain sizes. At larger problem sizes this issue, whilst still present, is ameliorated.
We then explored performance on a NVIDIA V100 GPU accelerator which is reported in fig. 10(b). On the GPU xDSL considerably outperforms PSyclone with the NVIDIA compiler for the PW advection benchmark and this is due to how data allocation is handled. PSyclone uses the managed memory option, whereas the xDSL GPU lowering detects that explicit memory allocation on the device will be more performant and handles this directly. Consequently, when the PSyclone compiled PW advection benchmarks were executed we found that there were a large number of unified memory GPU page faults which do not occur with xDSL.
The performance pattern for tracer advection on GPUs demonstrated in fig. 10(b) is similar to that of a single node illustrated in fig. 10(a), where xDSL lags PSyclone, especially for smaller problem sizes. This is also because of limitations in MLIR lowerings, this time from scf.parallel to the GPU dialect, where MLIR invokes a synchronous kernel execution for each parallel loop and thus execution across the CPU and GPU blocks at the end of each kernel.
| Benchmark | Throughput (GPts/s) | Performance | |
| Initial | Optimized | improvement | |
| pw-8m | 100x | ||
| pw-33 | 165x | ||
| pw-134m | 175x | ||
| traadv-4m | 113x | ||
| traadv-32m | 214x | ||
Table 1 reports performance for both benchmarks on an Alveo U280 FPGA. We support FPGAs by the development of an MLIR HLS dialect and transformations (Rodriguez-Canal et al., 2023b) which lower from this to LLVM-IR and others which lower to the HLS dialect from the stencil dialect. LLVM-IR is then provided to the AMD Xilinx HLS LLVM backend which synthesizes the corresponding kernel (Rodriguez-Canal et al., 2023a). The Initial version represents the algorithm running on the FPGA unchanged from it’s Von Neumann based CPU design, whereas the optimized version has been transformed by the compiler into a form tuned for dataflow architectures. This transformation has involved breaking the constituent components of the algorithm into separate dataflow regions and the use of a 3D shift buffer (Brown, 2021) which is a bespoke cache that, when filled, enables all the current grid cell’s stencil values to be provided to the calculation each cycle but one value needs to be read from DDR external memory per cycle. Ultimately, this means that calculation loops are able to be pipelined and run without stalling, computing a grid cell each cycle.
These performance numbers demonstrate two important benefits of our flow, firstly that we provide the ability to target FPGAs from a single Fortran source code, adding a new capability to PSyclone. Secondly, as can be seen by the performance improvement between optimized and initial versions in table 1, working with a high-level stencil description of the problem enables transformations to undertake automatic tuning to a dataflow architecture. Manual FPGA tuning is a very time-consuming step that requires considerable expertise (Brown, 2020), and so the ability for xDSL to automatically undertake this provides an important capability that has the potential to enhance vendor tools. Whilst the FPGA numbers reported in table 1 fall short of the NVIDIA V100 GPU performance, these are not only in agreement with the performance of tuned PW advection manually optimized for FPGAs (Brown, 2021), but it should also be highlighted that the MLIR GPU and CPU pipelines that we leverage are of production quality maturity and optimized by vendor teams.
Figure 11(a) reports the throughput for strong scaling of the PW advection benchmark on ARCHER2 up to 128 nodes (16384 cores). Unlike Devito, the PSyclone NEMO API we are using does not provide distributed memory capabilities, therefore, only xDSL-PSyclone results are presented here. The code was compiled with xDSL-PSyclone, and lowered using the stencil, dmp and mpi dialects, leveraging a 2D dmp decomposition strategy, which is commonplace in these types of model due to tight coupling in the vertical dimension. xDSL-PSyclone provides good strong scaling to eight nodes but then suffers from strong scaling effects due to the small global problem size. Likewise, fig. 11(b) reports throughput for the tracer advection benchmark where we observe similar behavior, with the 2D decomposition strategy limiting the strong scaling capability of xDSL-PSyclone.
7. Related Work
Numerous DSLs and compiler frameworks for HPC stencil computations exist. In contrast to building a siloed language and compiler infrastructure, our work differs from many previous attempts: we leverage the MLIR infrastructure to share maintenance support and interoperability with other community-driven IRs. High-level approaches to problem modeling using symbolic notation include the Python-based OpenSBLI (Jacobs et al., 2017), using Einstein notation and the C/C++-based OPS library (Reguly et al., 2018) underneath for automated parallelism. ExaStencils (Kuckuk and Köstler, 2016; Lengauer et al., 2020) has a multi-layered high-level symbolic DSL for FDs, targetting multi-CPU and multi-GPU. STELLA (Gysi et al., 2015) (and susbequently GridTools (Afanasyev et al., 2021)) uses a concise discretized mathematical syntax and helped port the COSMO model (Thaler et al., 2019) to NVIDIA GPUs.
Using a notation closer to modeling loops and arithmetic expressions, notable code-generation frameworks include Patus (Christen et al., 2011), the Pochoir stencil compiler (Tang et al., 2011), Physis (Maruyama et al., 2011) with C-based embedded DSL and Mint (Unat et al., 2011), a source-to-source translator for CUDA. MSC (Li et al., 2021) facilitates halo exchanges for DMP targeting many-core processor systems such as Sunway and Matrix. SDSLc (Rawat et al., 2015) targets CPUs and GPUs, Shift Calculus (Liao et al., 2015) differs by using an ontology language. Lift (Steuwer et al., 2015), based upon semantic-preserving rewrite rules to optimize stencils, StencilGen (Rawat et al., 2018) targets GPUs with more exotic optimizations, such as overlapped temporal blocking, and Artemis (Rawat et al., 2019) primarily targets GPUs. Other efforts to benefit from MLIR include domain-specific code generators using tensor-level abstractions (Essadki et al., 2023).
Targetting diverse problem domains in science other than FD stencils, for unstructured meshes, we encounter OP2 (Reguly et al., 2016; Mudalige et al., 2019), Liszt (DeVito et al., 2011), Paralab/Finch (Heisler et al., 2023), Firedrake (Ham et al., 2023), UFL (Alnæs, 2012) and FEniCS (Logg et al., 2012). SPIRAL (Puschel et al., 2005) targets Digital Signal Processing and Tensor Contraction Engine (Auer et al., 2006) for molecular physics. On the medical applications side, MOD2IR (Mitenkov et al., 2023) focuses on neuronal simulations, leveraging the LLVM toolchain to facilitate the optimization of NMODL (Hines and Carnevale, 2000). limpetMLIR (Thangamani et al., 2023) uses MLIR and targets electrophysiology models, and Stride (Cueto et al., 2022), built on top of Devito, focuses on ultra-sound tomography imaging. For tensor algebra, TACO (Kjolstad et al., 2017) offers a high-level DSL targeting a range of hardware accelerators. On the image processing side, Halide (Ragan-Kelley et al., 2013; Denniston et al., 2016) is a toolkit for generating HPC codes from functional specifications in image processing using programmer-specified scheduling transformations, the polyhedral-based PolyMage (Mullapudi et al., 2015) offers a high-level Python DSL and optimizing code generator and Forma (Ravishankar et al., 2015) also targetting GPUs and claims to provide an easier-to-handle DSL interface.
Focusing more on libraries and frameworks optimizing loop-dominated codes, Python-based Loo.py (Klöckner, 2014), generates code for OpenCL/CUDA, Chill (Chen et al., 2008) guides high-level loop transformations, YASK (Yount et al., 2016) specializing in data layout transformations to optimize for high-bandwidth Intel architectures. Numba (Lam et al., 2015) uses decorators for runtime-acceleration of Python code using LLVM as a backend. Notable annotation-based approaches include Tai-Chi (Hu et al., 2019) and DaCe (Ziogas et al., 2021).
Regarding frameworks to assist DSL developers in building DSLs for HPC code, AnyDSL (Leißa et al., 2018) provides composable DSL development using partial evaluation. Julia (Bezanson et al., 2017) combines dynamic types with the ability to generate HPC code. Python-based projects like Codon (Shajii et al., 2023), features static typing and compilation via LLVM, showcasing notable speedups compared to native Python routines and Mojo (Modular, [n. d.]) as a Pythonic frontend for MLIR, displays a significant performance boost over state-of-the-art tools. Finally, the MSF’s (Coullon et al., 2019) approach differs in addressing the sharing of common abstractions between stencil frameworks, providing a stencil-DSL operating in an ahead-of-time non-embedded fashion.
Describing and optimizing communication patterns for DMP/MPI scientific software is a focus of many libraries and compilers (Zhao et al., 2018; Wang and Chandramowlishwaran, 2020; Zhao et al., 2021; Castillo et al., 2019). Halide (Denniston et al., 2016), Exaslang (Kuckuk and Köstler, 2016) and PETSc (Zhang et al., 2022) facilitate large-scale computational science with libraries supporting efficient execution at all parallelism levels (SIMD, SMP, DMP). Finally, ParallelStencil.jl (Omlin et al., 2022) is a high-level Julia approach using bindings for MPI MPI.jl (Byrne et al., 2021).
8. Conclusions and Future Work
By merging the backend compilation stacks of two stencil-DSLs focusing on FD stencils, PSyclone and Devito, we presented a prototype to support these important domain-specific abstractions and their optimizations that can be shared across DSLs. As a result, their developer communities can benefit from delivering peak performance for key applications reasoning about FD stencils, such as seismic and medical imaging and weather modeling, targeting CPUs, GPUs, and clusters thereof. To enable broad sharing of abstractions, we expressed fundamental HPC abstractions such as message passing and halo exchanges as SSA-based IRs for the MLIR compiler ecosystem. We demonstrated that these abstractions enable the generation of code employing shared and distributed parallelism exhibiting competitive single-node and strong scaling performance to state-of-the-art supercomputers. The MPI abstractions are generic and independent of the specific application domain, while the DMP abstractions are focused on halo exchanges specific to FD stencil computations. Both enable the sharing of optimizations and code generation strategies for message passing across stencil computations that target supercomputers. Our evaluation across stencil kernel workloads from Devito and PSyclone shows competitive and even improved performance while broadening the set of targets.
We hope that this work on HPC abstractions for MPI and distributed-memory code generation transcends the specific application domains of our DSLs and motivates further research on exploiting synergies enabled by collaborations across DSL and compiler frameworks in HPC. We aim to establish a cohesive ecosystem of HPC-oriented IRs within the compiler landscape, facilitating reasoning and HPC code generation. These IRs promote modularity and interoperability by abstracting away from domain-specific considerations, aiming to enable code generation for efficient and scalable codes across DSLs targeting problems other than stencils.
A limitation identified by our benchmarking has been in the existing MLIR lowerings from the scf dialect to the OpenMP and GPU dialects. Further work optimizing these would be beneficial, potentially requiring developing bespoke lowerings which are able to consider the wider context that an scf.parallel operation sits within. The ability to support one parallel region across all scf.parallel loops in OpenMP and finer-grained control over the synchronization of GPU kernels are important aspects that would likely deliver significant performance improvements for codes like the tracer advection benchmark. Further work includes enhancing the decomposition techniques to support advanced exchange schemes and DMP/MPI optimizations, such as diagonal communications in the cartesian grid topology and communication/computation overlap. The abstractions presented in this paper aim to be modular and extensible building blocks to facilitate these future developments.
Acknowledgements.
This work was supported by the Sponsor Engineering and Physical Sciences Research Council (EPSRC) https://www.ukri.org/councils/epsrc/ grants Grant #EP/W007789/1 and Grant #EP/W007940/1. The authors would like to thank the shepherd and the reviewers for their comments and allocated time towards improving this work. The authors thank the xDSL, Devito, and PsyClone communities for their comments and discussions.References
- (1)
- Abadi et al. (2015) Martín Abadi, Ashish Agarwal, Paul Barham, Eugene Brevdo, Zhifeng Chen, Craig Citro, Greg S. Corrado, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Ian Goodfellow, Andrew Harp, Geoffrey Irving, Michael Isard, Yangqing Jia, Rafal Jozefowicz, Lukasz Kaiser, Manjunath Kudlur, Josh Levenberg, Dandelion Mané, Rajat Monga, Sherry Moore, Derek Murray, Chris Olah, Mike Schuster, Jonathon Shlens, Benoit Steiner, Ilya Sutskever, Kunal Talwar, Paul Tucker, Vincent Vanhoucke, Vijay Vasudevan, Fernanda Viégas, Oriol Vinyals, Pete Warden, Martin Wattenberg, Martin Wicke, Yuan Yu, and Xiaoqiang Zheng. 2015. TensorFlow: Large-Scale Machine Learning on Heterogeneous Systems. https://www.tensorflow.org/ Software available from tensorflow.org.
- Abadi et al. (2016) Martín Abadi, Paul Barham, Jianmin Chen, Zhifeng Chen, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Geoffrey Irving, Michael Isard, Manjunath Kudlur, Josh Levenberg, Rajat Monga, Sherry Moore, Derek G. Murray, Benoit Steiner, Paul Tucker, Vijay Vasudevan, Pete Warden, Martin Wicke, Yuan Yu, and Xiaoqiang Zheng. 2016. TensorFlow: a system for large-scale machine learning. In Proceedings of the 12th USENIX Conference on Operating Systems Design and Implementation (Savannah, GA, USA) (OSDI’16). USENIX Association, USA, 265–283. https://dl.acm.org/doi/10.5555/3026877.3026899
- Adams et al. (2019) S. V. Adams, R. W. Ford, M. Hambley, J. M. Hobson, I. Kavčič, C. M. Maynard, T. Melvin, E. H. Müller, S. Mullerworth, A. R. Porter, M. Rezny, B. J. Shipway, and R. Wong. 2019. LFRic: Meeting the challenges of scalability and performance portability in Weather and Climate models. J. Parallel and Distrib. Comput. 132 (2019), 383–396. https://doi.org/10.1016/j.jpdc.2019.02.007
- Afanasyev et al. (2021) Anton Afanasyev, Mauro Bianco, Lukas Mosimann, Carlos Osuna, Felix Thaler, Hannes Vogt, Oliver Fuhrer, Joost VandeVondele, and Thomas C. Schulthess. 2021. GridTools: a framework for portable weather and climate applications. SoftwareX 15 (2021), 100707. https://doi.org/10.1016/j.softx.2021.100707
- Alnæs (2012) Martin Sandve Alnæs. 2012. UFL: a finite element form language. Springer Berlin Heidelberg, Berlin, Heidelberg, 303–338. https://doi.org/10.1007/978-3-642-23099-8_17
- Auer et al. (2006) Alexander A. Auer, Gerald Baumgartner, David E. Bernholdt, Alina Bibireata, Venkatesh Choppella, Daniel Cociorva, Xiaoyang Gao, Robert Harrison, Sriram Krishnamoorthy, Sandhya Krishnan, Chi-Chung Lam, Qingda Lu, Marcel Nooijen, Russell Pitzer, J. Ramanujam, P. Sadayappan, and Alexander Sibiryakov. 2006. Automatic code generation for many-body electronic structure methods: the tensor contraction engine. Molecular Physics 104, 2 (2006), 211–228. https://doi.org/10.1080/00268970500275780
- Bezanson et al. (2017) Jeff Bezanson, Alan Edelman, Stefan Karpinski, and Viral B. Shah. 2017. Julia: A fresh approach to numerical computing. SIAM Rev. 59, 1 (Sept. 2017), 65–98. https://doi.org/10.1137/141000671
- Bisbas et al. (2021) George Bisbas, Fabio Luporini, Mathias Louboutin, Rhodri Nelson, Gerard J. Gorman, and Paul H. J. Kelly. 2021. Temporal blocking of finite-difference stencil operators with sparse “off-the-grid” sources. In 2021 IEEE International Parallel and Distributed Processing Symposium (IPDPS). 497–506. https://doi.org/10.1109/IPDPS49936.2021.00058
- Bisbas et al. (2023) George Bisbas, Rhodri Nelson, Mathias Louboutin, Paul H. J. Kelly, Fabio Luporini, and Gerard Gorman. 2023. Automated MPI code generation for scalable finite-difference solvers. arXiv:2312.13094 [cs.DC]
- Brown (2020) Nick Brown. 2020. Exploring the acceleration of Nekbone on reconfigurable architectures. In 2020 IEEE/ACM International Workshop on Heterogeneous High-performance Reconfigurable Computing (H2RC). 19–28. https://doi.org/10.1109/H2RC51942.2020.00008
- Brown (2021) Nick Brown. 2021. Accelerating advection for atmospheric modelling on Xilinx and Intel FPGAs. In 2021 IEEE International Conference on Cluster Computing (CLUSTER). 767–774. https://doi.org/10.1109/Cluster48925.2021.00113
- Brown et al. (2015) Nick Brown, Michele Weiland, Adrian Hill, Ben Shipway, Chris Maynard, Thomas Allen, and Mike Rezny. 2015. A highly scalable Met Office NERC Cloud model. In Proceedings of the 3rd International Conference on Exascale Applications and Software (Edinburgh, UK) (EASC ’15). University of Edinburgh, GBR, 132–137.
- Byrne et al. (2021) Simon Byrne, Lucas C. Wilcox, and Valentin Churavy. 2021. MPI.jl: Julia bindings for the Message Passing Interface. Proceedings of the JuliaCon Conferences 1, 1 (2021), 68. https://doi.org/10.21105/jcon.00068
- Castillo et al. (2019) Emilio Castillo, Nikhil Jain, Marc Casas, Miquel Moreto, Martin Schulz, Ramon Beivide, Mateo Valero, and Abhinav Bhatele. 2019. Optimizing Computation-Communication Overlap in Asynchronous Task-Based Programs. In Proceedings of the ACM International Conference on Supercomputing (Phoenix, Arizona) (ICS ’19). Association for Computing Machinery, New York, NY, USA, 380–391. https://doi.org/10.1145/3330345.3330379
- Chen et al. (2008) Chun Chen, Jacqueline Chame, and Mary Hall. 2008. CHiLL: A framework for composing high-level loop transformations. Technical Report 08-897. University of Southern California. https://citeseerx.ist.psu.edu/document?doi=6a4620589f63f3385707d2d590f7b7dc8ee4d74f
- Christen et al. (2011) Matthias Christen, Olaf Schenk, and Helmar Burkhart. 2011. PATUS: a Code Generation and Autotuning Framework for Parallel Iterative Stencil Computations on Modern Microarchitectures. In Proceedings of the 2011 IEEE International Parallel & Distributed Processing Symposium (IPDPS ’11). IEEE Computer Society, Washington, DC, USA, 676–687. https://doi.org/10.1109/IPDPS.2011.70
- Coullon et al. (2019) Hélène Coullon, Julien Bigot, and Christian Pérez. 2019. Extensibility and composability of a multi-stencil domain specific framework. International Journal of Parallel Programming 47 (2019), 1046–1085. https://doi.org/10.1007/s10766-017-0539-5
- Cueto et al. (2022) Carlos Cueto, Oscar Bates, George Strong, Javier Cudeiro, Fabio Luporini, Òscar Calderón Agudo, Gerard Gorman, Lluis Guasch, and Meng-Xing Tang. 2022. Stride: A flexible software platform for high-performance ultrasound computed tomography. Computer Methods and Programs in Biomedicine 221 (2022), 106855. https://doi.org/10.1016/j.cmpb.2022.106855
- Cytron et al. (1991) Ron Cytron, Jeanne Ferrante, Barry K. Rosen, Mark N. Wegman, and F. Kenneth Zadeck. 1991. Efficiently computing static single assignment form and the control dependence graph. ACM Trans. Program. Lang. Syst. 13, 4 (oct 1991), 451–490. https://doi.org/10.1145/115372.115320
- Denniston et al. (2016) Tyler Denniston, Shoaib Kamil, and Saman Amarasinghe. 2016. Distributed Halide. In Proceedings of the 21st ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming (Barcelona, Spain) (PPoPP ’16). Association for Computing Machinery, New York, NY, USA, Article 5, 12 pages. https://doi.org/10.1145/2851141.2851157
- DeVito et al. (2011) Zachary DeVito, Niels Joubert, Francisco Palacios, Stephen Oakley, Montserrat Medina, Mike Barrientos, Erich Elsen, Frank Ham, Alex Aiken, Karthik Duraisamy, Eric Darve, Juan Alonso, and Pat Hanrahan. 2011. Liszt: A domain specific language for building portable mesh-based PDE solvers. In SC ’11: Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis. 1–12. https://doi.org/10.1145/2063384.2063396
- Eldridge et al. (2021) Schuyler Eldridge, Prithayan Barua, Aliaksei Chapyzhenka, Adam Izraelevitz, Jack Koenig, Chris Lattner, Andrew Lenharth, George Leontiev, Fabian Schuiki, Ram Sunder, Andrew Young, and Richard Xia. 2021. MLIR as hardware compiler infrastructure. In Workshop on Open-Source EDA Technology (WOSET). https://woset-workshop.github.io/PDFs/2021/a06.pdf
- Essadki et al. (2023) Mohamed Essadki, Bertrand Michel, Bruno Maugars, Oleksandr Zinenko, Nicolas Vasilache, and Albert Cohen. 2023. Code Generation for In-Place Stencils. In Proceedings of the 21st ACM/IEEE International Symposium on Code Generation and Optimization (Montréal, QC, Canada) (CGO 2023). Association for Computing Machinery, New York, NY, USA, 2–13. https://doi.org/10.1145/3579990.3580006
- Fehr et al. (2023) Mathieu Fehr, Michel Weber, Christian Ulmann, Alexandre Lopoukhine, Martin Lücke, Théo Degioanni, Michel Steuwer, and Tobias Grosser. 2023. Sidekick compilation with xDSL. arXiv:2311.07422 [cs.PL]
- Govindarajan and Moses (2020) Sanath Govindarajan and William S. Moses. 2020. SyFER-MLIR: Integrating Fully Homomorphic Encryption Into the MLIR Compiler Framework. https://math.mit.edu/research/highschool/primes/materials/2020/Govindarajan-Moses.pdf
- Gysi et al. (2021) Tobias Gysi, Christoph Müller, Oleksandr Zinenko, Stephan Herhut, Eddie Davis, Tobias Wicky, Oliver Fuhrer, Torsten Hoefler, and Tobias Grosser. 2021. Domain-Specific Multi-Level IR Rewriting for GPU: The Open Earth Compiler for GPU-accelerated Climate Simulation. ACM Trans. Archit. Code Optim. 18, 4 (2021), 51:1–51:23. https://doi.org/10.1145/3469030
- Gysi et al. (2015) Tobias Gysi, Carlos Osuna, Oliver Fuhrer, Mauro Bianco, and Thomas C. Schulthess. 2015. STELLA: a domain-specific tool for structured grid methods in weather and climate models. In Proceedings of the International Conference for High Performance Computing, Networking, Storage and Analysis, SC 2015, Austin, TX, USA, November 15-20, 2015, Jackie Kern and Jeffrey S. Vetter (Eds.). ACM, New York, NY, USA, 41:1–41:12. https://doi.org/10.1145/2807591.2807627
- Ham et al. (2023) David A. Ham, Paul H. J. Kelly, Lawrence Mitchell, Colin J. Cotter, Robert C. Kirby, Koki Sagiyama, Nacime Bouziani, Sophia Vorderwuelbecke, Thomas J. Gregory, Jack Betteridge, Daniel R. Shapero, Reuben W. Nixon-Hill, Connor J. Ward, Patrick E. Farrell, Pablo D. Brubeck, India Marsden, Thomas H. Gibson, Miklós Homolya, Tianjiao Sun, Andrew T. T. McRae, Fabio Luporini, Alastair Gregory, Michael Lange, Simon W. Funke, Florian Rathgeber, Gheorghe-Teodor Bercea, and Graham R. Markall. 2023. Firedrake User Manual (first ed.). Imperial College London and University of Oxford and Baylor University and University of Washington. https://doi.org/10.25561/104839
- Hammond et al. (2023) Jeff Hammond, Lisandro Dalcin, Erik Schnetter, Marc Pérache, Jean-Baptiste Besnard, Jed Brown, Gonzalo Brito Gadeschi, Simon Byrne, Joseph Schuchart, and Hui Zhou. 2023. MPI Application Binary Interface Standardization. In Proceedings of the 30th European MPI Users’ Group Meeting (<conf-loc>, <city>Bristol</city>, <country>United Kingdom</country>, </conf-loc>) (EuroMPI ’23). Association for Computing Machinery, New York, NY, USA, Article 1, 12 pages. https://doi.org/10.1145/3615318.3615319
- Heisler et al. (2023) Eric Heisler, Aadesh Deshmukh, Sandip Mazumder, Ponnuswamy Sadayappan, and Hari Sundar. 2023. Multi-discretization domain specific language and code generation for differential equations. Journal of Computational Science 68 (2023), 101981. https://doi.org/10.1016/j.jocs.2023.101981
- Henretty et al. (2011) Tom Henretty, Kevin Stock, Louis-Noël Pouchet, Franz Franchetti, J. Ramanujam, and P. Sadayappan. 2011. Data Layout Transformation for Stencil Computations on Short-Vector SIMD Architectures. In Compiler Construction, Jens Knoop (Ed.). Springer Berlin Heidelberg, Berlin, Heidelberg, 225–245. https://doi.org/10.1007/978-3-642-19861-8_13
- Hines and Carnevale (2000) M. L. Hines and N. T. Carnevale. 2000. Expanding NEURON’s Repertoire of Mechanisms with NMODL. Neural Computation 12, 5 (05 2000), 995–1007. https://doi.org/10.1162/089976600300015475
- Hu et al. (2019) Yuanming Hu, Tzu-Mao Li, Luke Anderson, Jonathan Ragan-Kelley, and Frédo Durand. 2019. Taichi: A Language for High-Performance Computation on Spatially Sparse Data Structures. ACM Trans. Graph. 38, 6, Article 201 (2019), 16 pages. https://doi.org/10.1145/3355089.3356506
- Jacobs et al. (2017) Christian T. Jacobs, Satya P. Jammy, and Neil D. Sandham. 2017. OpenSBLI: A framework for the automated derivation and parallel execution of finite difference solvers on a range of computer architectures. Journal of Computational Science 18 (2017), 12–23. https://doi.org/10.1016/j.jocs.2016.11.001
- Kjolstad et al. (2017) Fredrik Kjolstad, Shoaib Kamil, Stephen Chou, David Lugato, and Saman Amarasinghe. 2017. The tensor algebra compiler. Proc. ACM Program. Lang. 1, OOPSLA, Article 77 (2017), 29 pages. https://doi.org/10.1145/3133901
- Klöckner (2014) Andreas Klöckner. 2014. Loo.py: transformation-based code generation for GPUs and CPUs. In Proceedings of ACM SIGPLAN International Workshop on Libraries, Languages, and Compilers for Array Programming (Edinburgh, United Kingdom) (ARRAY’14). Association for Computing Machinery, New York, NY, USA, 82–87. https://doi.org/10.1145/2627373.2627387
- Kuckuk and Köstler (2016) Sebastian Kuckuk and Harald Köstler. 2016. Automatic Generation of Massively Parallel Codes from ExaSlang. Computation 4, 3 (2016). https://doi.org/10.3390/computation4030027
- Lam et al. (2015) Siu Kwan Lam, Antoine Pitrou, and Stanley Seibert. 2015. Numba: a LLVM-based Python JIT compiler. In Proceedings of the Second Workshop on the LLVM Compiler Infrastructure in HPC (Austin, Texas) (LLVM ’15). Association for Computing Machinery, New York, NY, USA, Article 7, 6 pages. https://doi.org/10.1145/2833157.2833162
- Lattner et al. (2021) Chris Lattner, Mehdi Amini, Uday Bondhugula, Albert Cohen, Andy Davis, Jacques Pienaar, River Riddle, Tatiana Shpeisman, Nicolas Vasilache, and Oleksandr Zinenko. 2021. MLIR: Scaling Compiler Infrastructure for Domain Specific Computation. In 2021 IEEE/ACM International Symposium on Code Generation and Optimization (CGO). IEEE, 2–14. https://doi.org/10.1109/CGO51591.2021.9370308
- Leißa et al. (2018) Roland Leißa, Klaas Boesche, Sebastian Hack, Arsène Pérard-Gayot, Richard Membarth, Philipp Slusallek, André Müller, and Bertil Schmidt. 2018. AnyDSL: a partial evaluation framework for programming high-performance libraries. Proc. ACM Program. Lang. 2, OOPSLA (2018), 119:1–119:30. https://doi.org/10.1145/3276489
- Lengauer et al. (2020) Christian Lengauer, Sven Apel, Matthias Bolten, Shigeru Chiba, Ulrich Rüde, Jürgen Teich, Armin Größlinger, Frank Hannig, Harald Köstler, Lisa Claus, Alexander Grebhahn, Stefan Groth, Stefan Kronawitter, Sebastian Kuckuk, Hannah Rittich, Christian Schmitt, and Jonas Schmitt. 2020. ExaStencils: Advanced Multigrid Solver Generation. In Software for Exascale Computing - SPPEXA 2016-2019, Hans-Joachim Bungartz, Severin Reiz, Benjamin Uekermann, Philipp Neumann, and Wolfgang E. Nagel (Eds.). Springer International Publishing, Cham, 405–452. https://doi.org/10.1007/978-3-030-47956-5_14
- Li et al. (2024) Mingzhen Li, Bangduo Chen, Hailong Yang, Zhongzhi Luan, and Depei Qian. 2024. Building a domain-specific compiler for emerging processors with a reusable approach. In Science China Information Sciences, Vol. 67. https://doi.org/10.1007/s11432-022-3727-6
- Li et al. (2021) Mingzhen Li, Yi Liu, Hailong Yang, Yongmin Hu, Qingxiao Sun, Bangduo Chen, Xin You, Xiaoyan Liu, Zhongzhi Luan, and Depei Qian. 2021. Automatic Code Generation and Optimization of Large-Scale Stencil Computation on Many-Core Processors. In Proceedings of the 50th International Conference on Parallel Processing (Lemont, IL, USA) (ICPP ’21). ACM, New York, NY, USA, Article 34, 12 pages. https://doi.org/10.1145/3472456.3473517
- Liao et al. (2015) Chunhua Liao, Pei-Hung Lin, Daniel J. Quinlan, Yue Zhao, and Xipeng Shen. 2015. Enhancing Domain Specific Language Implementations through Ontology. In Proceedings of the 5th International Workshop on Domain-Specific Languages and High-Level Frameworks for High Performance Computing (Austin, Texas) (WOLFHPC ’15). ACM, New York, NY, USA, Article 3, 9 pages. https://doi.org/10.1145/2830018.2830022
- Logg et al. (2012) A. Logg, K.A. Mardal, and G. Wells (Eds.). 2012. Automated Solution of Differential Equations by the Finite Element Method: The FEniCS Book. Springer Berlin Heidelberg. https://doi.org/10.1007/978-3-642-23099-8
- Louboutin et al. (2019) M. Louboutin, M. Lange, F. Luporini, N. Kukreja, P. A. Witte, F. J. Herrmann, P. Velesko, and G. J. Gorman. 2019. Devito (v3.1.0): an embedded domain-specific language for finite differences and geophysical exploration. Geoscientific Model Development 12, 3 (2019), 1165–1187. https://doi.org/10.5194/gmd-12-1165-2019
- Luporini et al. (2020) Fabio Luporini, Mathias Louboutin, Michael Lange, Navjot Kukreja, Philipp Witte, Jan Hückelheim, Charles Yount, Paul H. J. Kelly, Felix J. Herrmann, and Gerard J. Gorman. 2020. Architecture and Performance of Devito, a System for Automated Stencil Computation. ACM Trans. Math. Softw. 46, 1, Article 6 (apr 2020), 28 pages. https://doi.org/10.1145/3374916
- Majumder and Bondhugula (2024) Kingshuk Majumder and Uday Bondhugula. 2024. HIR: An MLIR-based Intermediate Representation for Hardware Accelerator Description. , 13 pages. https://doi.org/10.1145/3623278.3624767
- Maruyama et al. (2011) Naoya Maruyama, Tatsuo Nomura, Kento Sato, and Satoshi Matsuoka. 2011. Physis: An Implicitly Parallel Programming Model for Stencil Computations on Large-Scale GPU-Accelerated Supercomputers. In Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis (Seattle, Washington) (SC ’11). ACM, New York, NY, USA, Article 11, 12 pages. https://doi.org/10.1145/2063384.2063398
- Melvin et al. (2017) Thomas Melvin, Steve Mullerworth, Rupert Ford, Chris Maynard, and Mike Hobson. 2017. LFRic: Building a new Unified Model. In EGU General Assembly Conference Abstracts. 13021.
- Meurer et al. (2017) Aaron Meurer, Christopher P. Smith, Mateusz Paprocki, Ondřej Čertík, Sergey B. Kirpichev, Matthew Rocklin, Amit Kumar, Sergiu Ivanov, Jason K. Moore, Sartaj Singh, Thilina Rathnayake, Sean Vig, Brian E. Granger, Richard P. Muller, Francesco Bonazzi, Harsh Gupta, Shivam Vats, Fredrik Johansson, Fabian Pedregosa, Matthew J. Curry, Andy R. Terrel, Štěpán Roučka, Ashutosh Saboo, Isuru Fernando, Sumith Kulal, Robert Cimrman, and Anthony Scopatz. 2017. SymPy: symbolic computing in Python. PeerJ Computer Science 3 (Jan. 2017), e103. https://doi.org/10.7717/peerj-cs.103
- Mitenkov et al. (2023) George Mitenkov, Ioannis Magkanaris, Omar Awile, Pramod Kumbhar, Felix Schürmann, and Alastair F. Donaldson. 2023. MOD2IR: High-Performance Code Generation for a Biophysically Detailed Neuronal Simulation DSL. In Proceedings of the 32nd ACM SIGPLAN International Conference on Compiler Construction (Montréal, QC, Canada) (CC 2023). ACM, New York, NY, USA, 203–215. https://doi.org/10.1145/3578360.3580268
- Modular ([n. d.]) Modular. [n. d.]. Mojo: Programming language for all of AI. https://www.modular.com/mojo
- Mudalige et al. (2019) G. R. Mudalige, I. Z. Reguly, S. P. Jammy, C. T. Jacobs, M. B. Giles, and N. D. Sandham. 2019. Large-scale performance of a DSL-based multi-block structured-mesh application for Direct Numerical Simulation. J. Parallel and Distrib. Comput. 131 (2019), 130–146. https://doi.org/10.1016/j.jpdc.2019.04.019
- Mullapudi et al. (2015) Ravi Teja Mullapudi, Vinay Vasista, and Uday Bondhugula. 2015. PolyMage: Automatic Optimization for Image Processing Pipelines. In Proceedings of the Twentieth International Conference on Architectural Support for Programming Languages and Operating Systems (Istanbul, Turkey) (ASPLOS ’15). Association for Computing Machinery, New York, NY, USA, 429–443. https://doi.org/10.1145/2694344.2694364
- Omlin et al. (2022) Samuel Omlin, Ludovic Räss, and Ivan Utkin. 2022. Distributed Parallelization of xPU Stencil Computations in Julia. arXiv:2211.15716 [cs.DC]
- Paszke et al. (2019) Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban Desmaison, Andreas Köpf, Edward Yang, Zach DeVito, Martin Raison, Alykhan Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu Fang, Junjie Bai, and Soumith Chintala. 2019. PyTorch: an imperative style, high-performance deep learning library. In Proceedings of the 33rd International Conference on Neural Information Processing Systems. Curran Associates Inc., Red Hook, NY, USA. https://dl.acm.org/doi/10.5555/3454287.3455008
- Piacsek and Williams (1970) Steve A Piacsek and Gareth P Williams. 1970. Conservation properties of convection difference schemes. J. Comput. Phys. 6, 3 (1970), 392–405. https://doi.org/10.1016/0021-9991(70)90038-0
- Porter et al. (2018) A. R. Porter, J. Appleyard, M. Ashworth, R. W. Ford, J. Holt, H. Liu, and G. D. Riley. 2018. Portable multi- and many-core performance for finite-difference or finite-element codes – application to the free-surface component of NEMO (NEMOLite2D 1.0). Geoscientific Model Development 11, 8 (2018), 3447–3464. https://doi.org/10.5194/gmd-11-3447-2018
- Puschel et al. (2005) M. Puschel, J. M. F. Moura, J. R. Johnson, D. Padua, M. M. Veloso, B. W. Singer, Jianxin Xiong, F. Franchetti, A. Gacic, Y. Voronenko, K. Chen, R. W. Johnson, and N. Rizzolo. 2005. SPIRAL: Code Generation for DSP Transforms. Proc. IEEE 93, 2 (2005), 232–275. https://doi.org/10.1109/JPROC.2004.840306
- Ragan-Kelley et al. (2013) Jonathan Ragan-Kelley, Connelly Barnes, Andrew Adams, Sylvain Paris, Frédo Durand, and Saman P. Amarasinghe. 2013. Halide: a language and compiler for optimizing parallelism, locality, and recomputation in image processing pipelines. In ACM SIGPLAN Conference on Programming Language Design and Implementation, PLDI ’13, Seattle, WA, USA, June 16-19, 2013, Hans-Juergen Boehm and Cormac Flanagan (Eds.). ACM, New York, NY, USA, 519–530. https://doi.org/10.1145/2491956.2462176
- Ravishankar et al. (2015) Mahesh Ravishankar, Justin Holewinski, and Vinod Grover. 2015. Forma: A DSL for Image Processing Applications to Target GPUs and Multi-Core CPUs. In Proceedings of the 8th Workshop on General Purpose Processing Using GPUs (San Francisco, CA, USA) (GPGPU-8). Association for Computing Machinery, New York, NY, USA, 109–120. https://doi.org/10.1145/2716282.2716290
- Rawat et al. (2015) Prashant Rawat, Martin Kong, Tom Henretty, Justin Holewinski, Kevin Stock, Louis-Noël Pouchet, J. Ramanujam, Atanas Rountev, and P. Sadayappan. 2015. SDSLc: A Multi-Target Domain-Specific Compiler for Stencil Computations. In Proceedings of the 5th International Workshop on Domain-Specific Languages and High-Level Frameworks for High Performance Computing (Austin, Texas) (WOLFHPC ’15). ACM, New York, NY, USA, Article 6, 10 pages. https://doi.org/10.1145/2830018.2830025
- Rawat et al. (2018) Prashant Singh Rawat, Miheer Vaidya, Aravind Sukumaran-Rajam, Mahesh Ravishankar, Vinod Grover, Atanas Rountev, Louis Noel Pouchet, and P. Sadayappan. 2018. Domain-Specific Optimization and Generation of High-Performance GPU Code for Stencil Computations. Proc. IEEE 106, 11 (2018), 1902–1920. https://doi.org/10.1109/JPROC.2018.2862896
- Rawat et al. (2019) Prashant Singh Rawat, Miheer Vaidya, Aravind Sukumaran-Rajam, Atanas Rountev, Louis-Noël Pouchet, and P. Sadayappan. 2019. On Optimizing Complex Stencils on GPUs. In 2019 IEEE International Parallel and Distributed Processing Symposium (IPDPS). IEEE, 641–652. https://doi.org/10.1109/IPDPS.2019.00073
- Reguly et al. (2016) István Z. Reguly, Gihan R. Mudalige, Carlo Bertolli, Michael B. Giles, Adam Betts, Paul H. J. Kelly, and David Radford. 2016. Acceleration of a Full-Scale Industrial CFD Application with OP2. IEEE Transactions on Parallel and Distributed Systems 27, 5 (2016), 1265–1278. https://doi.org/10.1109/TPDS.2015.2453972
- Reguly et al. (2018) István Z. Reguly, Gihan R. Mudalige, and Michael B. Giles. 2018. Loop Tiling in Large-Scale Stencil Codes at Run-Time with OPS. IEEE Trans. Parallel Distributed Syst. 29, 4 (2018), 873–886. https://doi.org/10.1109/TPDS.2017.2778161
- Rodriguez-Canal et al. (2023a) G. Rodriguez-Canal, N. Brown, T. Dykes, J. Jones, and U. Haus. 2023a. Fortran High-Level Synthesis: Reducing the Barriers to Accelerating HPC Codes on FPGAs. In 2023 33rd International Conference on Field-Programmable Logic and Applications (FPL). IEEE Computer Society, Los Alamitos, CA, USA, 10–18. https://doi.org/10.1109/FPL60245.2023.00010
- Rodriguez-Canal et al. (2023b) Gabriel Rodriguez-Canal, Nick Brown, Maurice Jamieson, Emilien Bauer, Anton Lydike, and Tobias Grosser. 2023b. Stencil-HMLS: A multi-layered approach to the automatic optimisation of stencil codes on FPGA. In Proceedings of the SC ’23 Workshops of The International Conference on High Performance Computing, Network, Storage, and Analysis (<conf-loc>, <city>Denver</city>, <state>CO</state>, <country>USA</country>, </conf-loc>) (SC-W ’23). Association for Computing Machinery, New York, NY, USA, 556–565. https://doi.org/10.1145/3624062.3624543
- Science and Technology Facilities Council (2021) (STFC) Science and Technology Facilities Council (STFC). 2021. PSycloneBench: small benchmarks used to inform the development of the PSyclone Domain-Specific Compiler. https://github.com/stfc/PSycloneBench Accessed 2023-04-24.
- Shajii et al. (2023) Ariya Shajii, Gabriel Ramirez, Haris Smajlović, Jessica Ray, Bonnie Berger, Saman Amarasinghe, and Ibrahim Numanagić. 2023. Codon: A Compiler for High-Performance Pythonic Applications and DSLs. In Proceedings of the 32nd ACM SIGPLAN International Conference on Compiler Construction (Montréal, QC, Canada) (CC 2023). ACM, New York, NY, USA, 191–202. https://doi.org/10.1145/3578360.3580275
- Steuwer et al. (2015) Michel Steuwer, Christian Fensch, Sam Lindley, and Christophe Dubach. 2015. Generating Performance Portable Code Using Rewrite Rules: From High-Level Functional Expressions to High-Performance OpenCL Code. SIGPLAN Not. 50, 9 (aug 2015), 205–217. https://doi.org/10.1145/2858949.2784754
- Tang et al. (2011) Yuan Tang, Rezaul Alam Chowdhury, Bradley C. Kuszmaul, Chi-Keung Luk, and Charles E. Leiserson. 2011. The Pochoir Stencil Compiler. In Proceedings of the Twenty-third Annual ACM Symposium on Parallelism in Algorithms and Architectures (San Jose, California, USA) (SPAA ’11). ACM, New York, NY, USA, 117–128. https://doi.org/10.1145/1989493.1989508
- Thaler et al. (2019) Felix Thaler, Stefan Moosbrugger, Carlos Osuna, Mauro Bianco, Hannes Vogt, Anton Afanasyev, Lukas Mosimann, Oliver Fuhrer, Thomas C. Schulthess, and Torsten Hoefler. 2019. Porting the COSMO Weather Model to Manycore CPUs. In Proceedings of the Platform for Advanced Scientific Computing Conference (Zurich, Switzerland) (PASC ’19). Association for Computing Machinery, New York, NY, USA, Article 13, 11 pages. https://doi.org/10.1145/3324989.3325723
- Thangamani et al. (2023) Arun Thangamani, Tiago Trevisan, Vincent Loechner, Stéphane Genaud, and Bérenger Bramas. 2023. Lifting Code Generation of Cardiac Physiology Simulation to Novel Compiler Technology. In 21st ACM/IEEE International Symposium on Code Generation and Optimization (CGO’23). ACM, ACM, New York, NY, USA, 68–80. https://doi.org/10.1145/3579990.3580008
- Unat et al. (2011) Didem Unat, Xing Cai, and Scott B. Baden. 2011. Mint: Realizing CUDA Performance in 3D Stencil Methods with Annotated C. In Proceedings of the International Conference on Supercomputing (Tucson, Arizona, USA) (ICS ’11). ACM, New York, NY, USA, 214–224. https://doi.org/10.1145/1995896.1995932
- Wang and Chandramowlishwaran (2020) Hengjie Wang and Aparna Chandramowlishwaran. 2020. Pencil: A Pipelined Algorithm for Distributed Stencils. In SC20: International Conference for High Performance Computing, Networking, Storage and Analysis. 1–16. https://doi.org/10.1109/SC41405.2020.00089
- Yount et al. (2016) Charles Yount, Josh Tobin, Alexander Breuer, and Alejandro Duran. 2016. YASK - Yet Another Stencil Kernel: A Framework for HPC Stencil Code-Generation and Tuning. In Sixth International Workshop on Domain-Specific Languages and High-Level Frameworks for High Performance Computing, WOLFHPC@SC 2016, Salt Lake, UT, USA, November 14, 2016. IEEE Computer Society, 30–39. https://doi.org/10.1109/WOLFHPC.2016.08
- Zhang et al. (2022) Junchao Zhang, Jed Brown, Satish Balay, Jacob Faibussowitsch, Matthew Knepley, Oana Marin, Richard Tran Mills, Todd Munson, Barry F. Smith, and Stefano Zampini. 2022. The PetscSF Scalable Communication Layer. IEEE Transactions on Parallel and Distributed Systems 33, 4 (2022), 842–853. https://doi.org/10.1109/TPDS.2021.3084070
- Zhao et al. (2021) Tuowen Zhao, Mary Hall, Hans Johansen, and Samuel Williams. 2021. Improving Communication by Optimizing On-Node Data Movement with Data Layout. In Proceedings of the 26th ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming (Virtual Event, Republic of Korea) (PPoPP ’21). Association for Computing Machinery, New York, NY, USA, 304–317. https://doi.org/10.1145/3437801.3441598
- Zhao et al. (2018) Tuowen Zhao, Samuel Williams, Mary Hall, and Hans Johansen. 2018. Delivering Performance-Portable Stencil Computations on CPUs and GPUs Using Bricks. In 2018 IEEE/ACM International Workshop on Performance, Portability and Productivity in HPC (P3HPC). 59–70. https://doi.org/10.1109/P3HPC.2018.00009
- Ziogas et al. (2021) Alexandros Nikolaos Ziogas, Timo Schneider, Tal Ben-Nun, Alexandru Calotoiu, Tiziano De Matteis, Johannes de Fine Licht, Luca Lavarini, and Torsten Hoefler. 2021. Productivity, Portability, Performance: Data-Centric Python. In Proceedings of the International Conference for High Performance Computing, Networking, Storage and Analysis (St. Louis, Missouri) (SC ’21). Association for Computing Machinery, New York, NY, USA, Article 95, 13 pages. https://doi.org/10.1145/3458817.3476176